Status Report

ISS On-Orbit Status 30 Jan 2003

By SpaceRef Editor
January 30, 2003
Filed under , ,
ISS On-Orbit Status 30 Jan 2003

All ISS systems continue to function nominally, except as noted previously
or below.

Wake-up was at the regular 1:00am EST, followed by post-sleep activities
and breakfast (begin of sleep time: 4:30pm).

FE-1 Nikolai Budarin and CDR Kenneth Bowersox worked on the Progress
9P spacecraft, installing the docking mechanism (StM) of the cargo vehicle’s
SSVP docking and internal transfer system. [The SSVP on Progress (and
Soyuz) is of the "probe and cone" ( "classic") type,
to distinguish it from the structurally more robust hybrid ASP type used
on ISS for connecting Russian modules together. Besides acting as structural
and mechanical connection allowing passage of crew and cargo, it also
has connections for power, data (command and control), gases and fluids

Afterwards, Budarin deinstalled and removed no-longer-needed electronic
equipment from the Progress, viz., a local signal switch device (LKT/TA251)
of the BITS onboard telemetry system, with its ROM memory (PZU), for
return to Earth and recycling on a future Progress.

The crew continued stowing disposable cargo on Progress, while video-taping
the activities with the U.S. Sony PD-100 DVCAM. The recording was then
downlinked via Ku- and S-band at 7:10am.

Budarin conducted his regular daily maintenance checkup of the Russian
BIO-5 Rasteniya-2/Lada-2 ("Plants-2") plant growth experiment.
Later, he copied its accumulated photo/data files to a floppy disk for
transfer to the Russian Laptop 3 and subsequent downlink via Regul-Packet.

FE-2/SO Don Pettit terminated the regeneration cycle on the first EMU
Metox (metal oxide) CO2 filter in the Airlock’s bake-out oven, then started
it on the second canister.

Early in the morning, Pettit took carbon dioxide partial pressure (ppCO2)
readings of the cabin air in Lab and SM with the portable ACS CDMK (atmosphere
control system/CO2 monitoring kit). A second data take is scheduled for
shortly before sleep time. [ppCO2 levels have shown a gradual increase,
going above the lower limit of 5.3 mmHg over the last weekend. The crew
has been instructed to resume daily CDMK readings (morning and evening)
for the near-term, while specialists investigate the cause. The Service
Module (SM) CO2 sensor is reading ppCO2 levels well below the Lab MCA
(major constituent analyzer) and CDMK measurements (see Systems Status,

Air samples were collected by Pettit in the Lab, SM, and Node with the
MAS (microbial air sampler) kit for bacterial and fungal analysis, to
be performed on Days 2 and 5 of incubation in Petri dishes. The collected
samples were photographed by Bowersox for the record.

Sox then conducted surface sample swabs in the Lab, SM and Node with
the SSK (surface sample kit), for cultivation. The colony growth on the
sampling slides will be analyzed after two and five days of incubation.
The activities and SSK samples were photographed by Don for the record

Troubleshooting continues on the ER3 ARIS (Express rack 3/active rack
isolation system), after the crew was unable to perform the pushrod calibration
on 1/28. [Under investigation is a possible issue with the new software
load (file transfer from the ER3 memory unit {EMU} to the ARIS controller).
To protect the ARIS hardware, the ARIS alignment guides need to remain
installed until the problem is fixed. Some crew support may be requested
for tomorrow (Friday).]

Early in the morning, Don Pettit powered up the HRF (human research
facility) for checking out the FOOT (foot/ground reaction forces during
space flight) experiment’s montage file, which he then copied for downlink.
The HRF computer’s data files were subsequently downlinked on ground
command, and the HRF was deactivated again.

Pettit also worked on the EarthKAM (EK), changing its optics to the
180-mm lens. Unattended (ground-commanded) uplink and downlink of EK
payload control and data files through OCA took place throughout the
day. [The EK team, on behalf of the 29 participating middle schools across
the USA, Mexico and Japan, sent up sincere thanks for the crew’s support
in setting up EK on the OpsLAN (operations/local area network). So far,
the ground has commanded and received over 280 spectacular images of
the Earth, and two favorite ones were uplinked for the crew’s perusal.
EarthKAM images can be found at ].

Budarin completed the regular daily task of SOSh life support systems
maintenance, and performed the periodic routine inspection of the BRPK
air/condensate water separator in the SM, while Pettit took care of the
daily IMS database update file preparation.

A Lab RPCM (remote power controller module), which powers the VTR1 (video
tape recorder #1), has tripped open. VTR1 was declared "off limits" until
the problem is resolved. [Telemetry revealed no evidence ofa VTRovercurrent
that would have tripped the fuse. Although this could simply be a nuisance
trip, the signature is similar to a problem seen in the past, which eventually
led to a POR (power-on reset). This would require the ground to run a
specific nominal procedure, with the VTR off.]

Early in the morning, MCC-H and its support group in Moscow (HSG) performed
a BCC (backup control center) dry-run in test mode, with no involvement
of ISS crew or vehicle. [In test mode, HSG (as BCC) was to send a single
not-for-uplink PPCP (preplanned command packet) to MCC-M/TsUP, then
to cancel it after successful transmission. MCC-H was to play back
the ACT (American contingency telemetry) recorded during the Hurricane
Lili contingency operations last year (when MCC-H was shut down at
4:00am EDT on 10/2), with HSG confirming successful receipt. TsUP commanding
to the Russian segment (RS) via U.S. S-band, as well as MCC-H-to-TsUP
flow of status telemetry were also not affected.]

Another water venting from the Lab condensate collection tank is tentatively
scheduled for Friday next week (2/7), when tank contents are predicted
to reach the upper limit of 100 lbs. During the dump, the Lab window
shutter will be closed by the crew to protect it against contamination.
[According to applicable Flight Rule, the condensate tank should preferably
be off-loaded into a CWC (contingency water container) for processing/recycling
in theRS. Currently, however, US condensate can not be processed until
about April this year due to previous failures of the Russian SRV-K water
processor system. The excess water must therefore be vented overboard
(dual-stream, to cancel out propulsive reaction on the vehicle), in LVLH
(+XVV) attitude.]

Today’s targets for the CEO (crew earth observations program) were Mekong
River Delta (pass crossed the entire width of the delta from southwest
to northeast. Detailed nadir mapping pass requested [looking a touch
left for sun glint views]), Saigon, Vietnam (nadir pass; ESC [electronic
still camera]. The city lies inland about 20 km on the easternmost river
course of the delta), Industrialized SE Africa haze (pass paralleled
the Great Escarpment: looking left for aerosol loadings on the plateau
and right for a probably different airmass on the coastal plain. Smog
loadings should increase with distance north from the coast as ISS approached
the industrial heartland of the Witwatersrand, off-track left ~1.5. [Nadir
views of the knot of mountains that make up the Kingdom of Lesotho are
of great interest since major new dams have been built to supply Johannesburg
with water 250 miles to the north]), Delhi, India (looking right of track
for the city which lies on both banks of the upper Ganges [looking for
converging pattern of railroads/highways. Crew was also to shoot high
obliques, looking right, along the Himalayan Mt front, to document the
remnants of major fog banks that have dominated the Ganges Plain for
weeks), Mt Cameroon (Dynamic event. Unusual clear weather in the Bight
of Benin should make this major volcano visible [the volcano drops into
the sea from 13000 feet]. Looking right for recent lava flows as well
as a major collapse canyon on the west side of the volcano. There are
almost no handheld images of Mt Cameroon), West Africa dust (Dynamic
event. ISS track crossed over the southern dust front of a new outblow
of Saharan dust. This plume was centered over the Guineas, aligned SW.
If possible, crew was to try to document the source region[s] of this
plume. Since the plume lies further south than most, sources probably
lie in the zone of disturbed Sahelian soils), and Bamako, Mali (looking
slightly right of track on the Niger River. ESC requested).

CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 1:25pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
Elektron O2 generator is powered On (16 Amp mode). Vozdukh CO2 scrubber
is Off. U.S. CDRA CO2 scrubber is On. TCCS (trace contaminant control
subsystem) is operating. MCA (major constituents analyzer) is operating.
BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2
in Purify mode. RS air conditioner SKV-1 is On; SKV-2 is Off.

SM Working Compartment: Pressure (mmHg) — 753; temperature (deg C)
— 23.7; ppO2 (mmHg) — 160.7; ppCO2 (mmHg) — 2.5 (suspect).
SM Transfer Compartment: Pressure (mmHg) — 751; temperature (deg C)
— 18.8.
FGB Cabin: Pressure (mmHg) — 752; temperature (deg C) — 21.3.
Node: Pressure (mmHg) — 738.16; temperature (deg C) — 22.8 (shell);
ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
U.S. Lab: Pressure (mmHg) — 740.78; temperature (deg C) — 22.5; ppO2
(mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock (Equip. Lock): Pressure (mmHg) — 740.88; temperature (deg
C) — 19.0; shell heater temp (deg C) — 20.5, ppO2 (mmHg) — 164.1;
ppCO2 (mmHg) — 5.3.
PMA-1: Shell heater temp (deg C) — 23.7
PMA-2: Shell heater temp (deg C) — 18.5
(n/a = data not available)

Propulsion System (PS): Total propellant load available [SM(820) + FGB(2898)
+ Progress(0)] — 3718 kg (8197 lb) as of 1/30/03. (Capability: SM —
860 kg; FGB — 6120 kg).

Electrical Power Systems (EPS):
Both P6 channels fully operational. Beta Gimbal Assembly (BGA) 2B and
BGA 4B both in "blind" dual-angle mode (directed position).
SM batteries: Battery #7 is off line (failed); battery #2 is offline
(ROM/capacity restoration mode); all other batteries (6) are in "Partial
Charge" mode.
FGB batteries: Battery #2 is offline; all other batteries (5) are in "Partial
Charge" mode.
Plasma Contactor Unit PCU-1 and PCU-2 both in Standby mode.

Command & Data Handling Systems:
C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
GNC-1 MDM is prime; GNC-2 is Backup (new patches loaded on both).
EXT-1 is On (primary), EXT-2 is off.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-2 MDM is On (primary); PL-1 MDM is off (diagnostic
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational. Lane
1 is down (as of 11/14/02).
SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:
3 CMGs on-line (CMG-1 failed).
State vector — U.S. SIGI-1
Attitude — U.S. SIGI-1
Angular rates — U.S. RGA-1

Flight Attitude:
LVLH TEA (local vertical/local horizontal = "earth-fixed":
z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch:
-9.1 deg, roll: 0 deg]), CMG/TA (thruster assist) momentum management).
Solar Beta Angle: 16.2 deg (magnitude decreasing).

Communications & Tracking Systems:
FGB MDM-1 is powered Off; FGB MDM-2 is operational.
All other Russian communications & tracking systems are nominal.
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally, but VTR1 is off.
HCOR (high-rate communications outage recorder) is operating nominally.

SSRMS/Canadarm2 based at MBS PDGF2 with Keep Alive (KA) power on both
MBS: KA power on both strings.
MT: latched at WS4, with KA power.
POA: KA power on both strings.
RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit (as of this morning, 5:24am EST [= epoch]):
Mean altitude — 387.7 km
Apogee — 391.5 km
Perigee — 383.9 km
Period — 92.31 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0005622
Orbits per 24-hr. day — 15.60
Mean altitude loss in last 24 hours — 200 m
Solar Beta Angle — 16.2 deg (magnitude decreasing)
Revolutions since FGB/Zarya launch (Nov. ’98) — 23951

For more on ISS orbit and worldwide naked-eye visibility dates/times,

SpaceRef staff editor.